OpenStudy (anonymous):

let f be a linear function where f(2)=-5 and f(-3)=1. state the function f(x)

OpenStudy (zarkon):

\[f(x)=ax+b\] \[f(2)=2a+b=-5\] \[f(-3)=-3a+b=1\] solve this system of two equations and two unknowns

OpenStudy (zarkon):

or you have the points (2,-5) and (-3,1) find the slope and then the equation is \[y-y_1=m(x-x_1)\]